# MF-GLaM: A multifidelity stochastic emulator using generalized lambda models

### v1

@@ 4 Applications
$d_{\mathrm{WS}}\left(Y_{1},Y_{2}\right)=\left\|Q_{1}-Q_{2}\right\|_{2}=\sqrt{\int_{0}^{1}\left(Q_{1}(u)-Q_{2}(u)\right)^{2}\mathrm{\leavevmode\nobreak\ d}u}.$ (35)
The expectation in Eq. 34 is approximated using the sample average over a test set $\boldsymbol{{\mathcal{X}}}_{\text{test}}=\{\boldsymbol{x}_{\text{test}}^{(1)},...,\boldsymbol{x}_{\text{test}}^{(N_{\text{test}})}\}$ of size $N_{\text{test}}=1,\!000$, generated via Latin hypercube sampling (LHS) (McKay et al., 1979). In the first application, where the true HF model is itself a GLaM, the quantiles to compute the Wasserstein distance in Eq. 35 are obtained directly from Eq. 3. Furthermore, the variance in the denominator of Eq. 34 is estimated using the law of total variance, expressed as:
@@
This variance is estimated on $\boldsymbol{{\mathcal{X}}}_{\text{test}}$, where the conditional means $\mathbb{E}_{\boldsymbol{X}}[Y^{\text{H}}_{\boldsymbol{x}}]$ and variances $\text{Var}(Y^{\text{H}}_{\boldsymbol{x}})$ at each point are obtained from Eqs. 5 and 6. In contrast, for the final two applications, where neither the analytical form of the true HF response distribution $Y_{\boldsymbol{x}}^{\text{H}}$ nor its moments and quantiles are available, we perform $N_{\text{R}}=10^{4}$ replications (i.e., repeated evaluations) of the HF stochastic simulator at each $\boldsymbol{x}_{\text{test}}^{(i)}$. Then, the quantiles of each response distribution are approximated by empirical quantiles across the replications, while the overall variance $\text{Var}[Y_{\text{H}}]$ is estimated as the empirical variance of the aggregated responses across all $N_{\text{R}}\times N_{\text{test}}$ evaluations.
@@
Across all three examples, the candidate PCE degrees and q-norms for the polynomial expansion of each parameter in the HF-only, LF-only and MF- GLaMs, are specified as shown in Table 1. When $N_{\mathrm{H}}=100$, the HF and discrepancy degrees are restricted to mitigate the risk of overfitting. Finally, the implementation of our code for the MF-GLaM methodology and its associated applications is built upon the existing GLaM module of UQLab version 2.1.0 (Marelli and Sudret, 2014; Lüthen et al., 2024), a general-purpose uncertainty quantification software developed in Matlab.
@@ 4.1 Synthetic GLaM models
In this first example, both the high-fidelity and low-fidelity stochastic simulators are GLaMs defined on a four-dimensional input space. As the true response distributions exactly follow the GLaM form, this case provides the minimal-complexity setting needed to validate our multifidelity procedure in the case of perfect model response agreement, before proceeding to more complex, imperfectly represented models. The input variables $X_{1},X_{2},X_{3},X_{4}$ are assumed to follow independent uniform distributions over $[0,2]$, i.e. $X_{1},X_{2},X_{3},X_{4}\sim\mathcal{U}([0,2])$. The four parameters $\boldsymbol{\lambda}^{\text{H}}$ and $\boldsymbol{\lambda}^{\text{L}}$ of the HF and LF model, respectively, are represented using polynomial chaos expansions, with the corresponding nonzero coefficients listed in Table 2. Each multi-index defines a polynomial basis function, where a value $p$ at the $i$-th position of a given multi-index corresponds to $x_{i}$ raised to the power of $p$. For example, the multi-index 2100 corresponds to the product of two Legendre polynomials: one of order $2$ in $x_{1}$, and one of order $1$ in $x_{2}$.
@@
To illustrate how the HF and LF models differ across the input domain, Fig. 1 compares their conditional response distributions at three representative input points. In this figure, HF responses are shown as black dashed curves, while LF responses appear as gray dotted curves. The same plots also compare the predictions from two stochastic emulators: an MF-GLaM (blue curves) and a GLaM trained solely on HF data (red curves), the latter referred to as the “HF-only” GLaM. Results are shown for increasing HF training set sizes, ranging from $N_{\text{H}}=100$ up to $N_{\text{H}}=400$ data points, while the LF training set size remains fixed at $N_{\text{L}}=1,\!000$ data points. We notice that, for all three input points and across all HF training set sizes, the MF-GLaM consistently captures the conditional HF distribution more accurately than the HF-only GLaM. Even for $\boldsymbol{x}_{2}$, where the LF and HF response supports are disjoint, a quite uncommon scenario in practice, the MF-GLaM reproduces the HF conditional PDF more faithfully compared to relying solely on HF data. We notice that in this four-dimensional application, $N_{\text{H}}=100$ HF training data points are insufficient to achieve an acceptable approximation of the response PDF; however, the MF approach still yields improved predictions compared to the HF-only model. Once the HF training set size increases to $N_{\text{H}}=200$, the MF predictions become very close to the true HF PDF, and with $N_{\text{H}}=400$, the MF approximation is nearly exact. In contrast, the HF-only GLaM exhibits greater variability and struggles to reproduce the PDF shape in certain cases (particularly for inputs $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$).
@@
These observations are further supported by the global approximation performance comparison shown in Fig. 2, and its numerical summary, provided in Table 3. The box plot presents the performance comparison in terms of normalized Wasserstein distance $\epsilon_{\text{W}}$ of the MF-GLaM, the HF-only GLaM, and the LF-only GLaM, the latter serving as a baseline. Each box group corresponds to a different HF training set size, ranging from $N_{\text{H}}=100$ to $N_{\text{H}}=800$ points, with $N_{\text{r}}=25$ repetitions per scenario. Throughout these tests, we drew once $N_{\text{r}}$ LF training datasets of size $N_{\text{L}}=1,\!000$ each, which were reused across all $N_{\text{H}}$ values.
We observe that the MF-GLaM consistently outperforms the GLaM trained solely on the HF data and shows faster convergence to the true HF model, indicated by the rapid decrease in Wasserstein distance. Especially for $N_{\text{H}}=200$, the error is reduced by approximately one order of magnitude. The only exception occurs at $N_{\text{H}}=100$, where the median performance of the MF and HF-only emulators appears comparable. This can be attributed to the imposed limitation of the correction term $\lambda_{1}^{\delta,\text{PC}}$ to a polynomial degree of 1 at this training size, introduced to avoid overfitting. However, as shown in Table 2, the actual discrepancy between the HF and LF expansions of $\lambda_{1}$ involves higher-degree terms. Nonetheless, the MF-GLaM demonstrates reduced variability across repetitions, as indicated by the significantly smaller interquartile range of $\epsilon_{\text{W}}$ in Table 3, which suggests more robust predictive performance.
Figure 2: Synthetic GLaMs example. Comparison of the normalized Wasserstein distance $\epsilon_{\text{W}}$ for LF-only, HF-only, and MF GLaMs across increasing HF experimental design sizes $N_{\text{H}}$. The box plots correspond to $N_{\text{r}}=25$ repetitions of the procedure, and $\epsilon_{\text{W}}$ is computed using $1,\!000$ test points.
Table 3: Synthetic GLaMs example. Median (interquartile range) of the normalized Wasserstein distance $\epsilon_{\text{W}}$$\times 10^{-2}$, computed over $N_{\text{r}}=25$ repetitions, comparing HF-only GLaM and MF-GLaM models for four HF training set sizes. Baseline $\epsilon_{\text{W}}$, for the LF-only GLaM: $5.8\,(0.1)\times 10^{-2}$.
@@
Finally, we also compare how well the emulators reproduce the mean and variance of the HF stochastic model. Since these two statistical quantities are deterministic, we use a normalized mean-square error (NMSE) measure to assess their prediction accuracy. Specifically, for each point $\boldsymbol{x}^{(i)}$ in a test set of size $N_{\mathrm{t}}=10^{4}$, the NMSE is given by:
@@
and $\eta_{\text{H}}^{(i)}$ the true value of the mean or variance at $\boldsymbol{x}^{(i)}$ for the HF simulator, given by Eq. 5 and Eq. 6, respectively, since in this example our true HF simulator is a GLaM. Also, $\eta_{\text{GLaM}}^{(i)}$ is the corresponding mean or variance value predicted by the GLaM surrogate model.
LABEL:fig:06Synth_RMSE_mean_var shows the NMSE for the mean (LABEL:fig:06Synth_RMSE_mean) and variance (LABEL:fig:06Synth_RMSE_var) estimates of the stochastic emulators whose predictive distributions were evaluated in Fig. 2. The corresponding box plots align with those for the Wasserstein distance and further confirm that the MF-GLaM provides more accurate and robust predictions compared to its single-fidelity counterparts, also for the statistics of the model response.
@@ 4.2 Stochastic borehole function
The input vector $\boldsymbol{x}$ contains eight random variables, whose distributions are shown in Table 4.
@@
[…]
@@
Moreover, Lüthen et al. (2023) modified the borehole function in Eq. 38 into a stochastic simulator by treating the five input parameters, $\boldsymbol{\Omega}_{\text{H}}=(R,T_{u},T_{l},H_{l},L)$, as latent variables. We adopt this variation for our application as the HF stochastic simulator, which results in the following three-dimensional HF model: $f_{s,\text{H}}(r_{w},h_{u},k_{w})=f(r_{w},h_{u},k_{w};\omega_{\text{H}})$.
Additionally, we obtain the LF stochastic simulator from the deterministic LF model in Eq. 39 by considering six of its input parameters to be latent: $\boldsymbol{\Omega}_{\text{L}}=(K_{w},R,T_{u},T_{l},H_{l},L)$. This results in the following two-dimensional LF model: $f_{s,\text{L}}(r_{w},h_{u})=f_{\text{L}}(r_{w},h_{u};\omega_{\text{L}})$.
This example investigates the applicability of the multifidelity GLaM methodology in cases where the LF explicit input variables are a subset of the HF ones. In this example, the two-dimensional LF input is a subset of the three-dimensional HF input. As a result, whereas the two common input variables can be modeled using all the available HF and LF data in the multifidelity GLaM through both the LF and the discrepancy part in Eq. 19, the additional HF-only input variable and its interactions have to be captured exclusively through the discrepancy component of the MF-GLaM, relying solely on the HF data.
Fig. 4 compares the PDF approximation capabilities of an MF-GLaM and an HF-only GLaM for three random points. In the same plots, we see the true HF PDF (black dashed curve) as well as the true LF (gray dotted curve) PDF. We notice that the HF-LF relationship can vary across the input space. For example, at $\boldsymbol{x}_{3}$, the LF PDF closely matches the HF reference, whereas at $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$ the LF distribution is more noticeably shifted or differs in shape. The MF-GLaM (blue curve) successfully corrects for these discrepancies, consistently aligning with the true HF response, even when $N_{\text{H}}=100$. In contrast, the HF-only GLaM (red curve) tends to misplace the peak of the PDF or produce much longer tails, especially at smaller training sizes.
@@
Fig. 5 and its numerical summary in Table 5 compare the global approximation capabilities of the LF-only, HF-only, and MF-GLaMs in terms of the normalized Wasserstein distance $\epsilon_{\text{W}}$. The size of the HF training set ranges from $N_{\text{H}}=100$ up to $N_{\text{H}}=800$, while the LF training set size is fixed at $N_{\text{L}}=1,\!000$ data points. Overall, the MF-GLaM (blue boxes) consistently outperforms the HF-only GLaM (red boxes), especially when the HF training set is small. Particularly, with only $N_{\text{H}}=100$ high-fidelity samples, the MF-GLaM achieves accuracy comparable to that of the HF-only model trained with $N_{\text{H}}=400$. As $N_{\text{H}}$ increases to 800, the gap between the MF and HF-only approaches narrows; however, such a large high-fidelity to low-fidelity data size ratio is generally atypical for multi-fidelity scenarios.
Figure 5: Stochastic borehole example. Comparison of the normalized Wasserstein distance $\epsilon_{\text{W}}$ for LF-only, HF-only, and MF GLaMs across increasing HF experimental design sizes $N_{\text{H}}$. The box plots correspond to 25 repetitions of the procedure, and $\epsilon_{\text{W}}$ is computed using $1,\!000$ test points.
Table 5: Stochastic borehole example. Median (interquartile range) of the normalized Wasserstein distance $\epsilon_{\text{W}}$$\times 10^{-2}$, computed over $N_{\text{r}}=25$ repetitions, comparing HF-only GLaM and MF-GLaM models for four HF training set sizes. Baseline $\epsilon_{\text{W}}$, for the LF-only GLaM: $2.9\,(0.2)\times 10^{-2}$.

### v2

@@ 4 Applications
$d_{\mathrm{WS}}\left(Y_{1},Y_{2}\right)=\left\|Q_{1}-Q_{2}\right\|_{2}=\sqrt{\int_{0}^{1}\left(Q_{1}(u)-Q_{2}(u)\right)^{2}\mathrm{\penalty 10000\ d}u}.$ (35)
The expectation in Eq.˜34 is approximated using the sample average over a test set $\boldsymbol{{\mathcal{X}}}_{\text{test}}=\{\boldsymbol{x}_{\text{test}}^{(1)},...,\boldsymbol{x}_{\text{test}}^{(N_{\text{test}})}\}$ of size $N_{\text{test}}=1,\!000$, generated via Latin hypercube sampling (LHS) (McKay et al., 1979). In the first application, where the true HF model is itself a GLaM, the quantiles to compute the Wasserstein distance in Eq.˜35 are obtained directly from Eq.˜3. Furthermore, the variance in the denominator of Eq.˜34 is estimated using the law of total variance, expressed as:
@@
This variance is estimated on $\boldsymbol{{\mathcal{X}}}_{\text{test}}$, where the conditional means $\mathbb{E}_{\boldsymbol{X}}[Y^{\text{H}}_{\boldsymbol{x}}]$ and variances $\text{Var}(Y^{\text{H}}_{\boldsymbol{x}})$ at each point are obtained from Eqs.˜5 and 6. In contrast, for the final two applications, where neither the analytical form of the true HF response distribution $Y_{\boldsymbol{x}}^{\text{H}}$ nor its moments and quantiles are available, we perform $N_{\text{R}}=10^{4}$ replications (i.e., repeated evaluations) of the HF stochastic simulator at each $\boldsymbol{x}_{\text{test}}^{(i)}$. Then, the quantiles of each response distribution are approximated by empirical quantiles across the replications, while the overall variance $\text{Var}[Y_{\text{H}}]$ is estimated as the empirical variance of the aggregated responses across all $N_{\text{R}}\times N_{\text{test}}$ evaluations.
@@
Across all three examples, the candidate PCE degrees and q-norms for the polynomial expansion of each parameter in the HF-only, LF-only and MF- GLaMs, are specified as shown in Table˜1. When $N_{\mathrm{H}}=100$, the HF and discrepancy degrees are restricted to mitigate the risk of overfitting. Finally, the implementation of our code for the MF-GLaM methodology and its associated applications is built upon the existing GLaM module of UQLab version 2.1.0 (Marelli and Sudret, 2014; Lüthen et al., 2024), a general-purpose uncertainty quantification software developed in Matlab.
@@ 4.1 Synthetic GLaM models
In this first example, both the high‑fidelity and low‑fidelity stochastic simulators are GLaMs defined on a four‑dimensional input space. As the true response distributions exactly follow the GLaM form, this case provides the minimal‑complexity setting needed to validate our multifidelity procedure in the case of perfect model response agreement, before proceeding to more complex, imperfectly represented models. The input variables $X_{1},X_{2},X_{3},X_{4}$ are assumed to follow independent uniform distributions over $[0,2]$, i.e. $X_{1},X_{2},X_{3},X_{4}\sim\mathcal{U}([0,2])$. The four parameters $\boldsymbol{\lambda}^{\text{H}}$ and $\boldsymbol{\lambda}^{\text{L}}$ of the HF and LF model, respectively, are represented using polynomial chaos expansions, with the corresponding nonzero coefficients listed in Table˜2. Each multi-index defines a polynomial basis function, where a value $p$ at the $i$-th position of a given multi-index corresponds to $x_{i}$ raised to the power of $p$. For example, the multi-index 2100 corresponds to the product of two Legendre polynomials: one of order $2$ in $x_{1}$, and one of order $1$ in $x_{2}$.
@@
To illustrate how the HF and LF models differ across the input domain, Fig.˜1 compares their conditional response distributions at three representative input points. In this figure, HF responses are shown as black dashed curves, while LF responses appear as gray dotted curves. The same plots also compare the predictions from two stochastic emulators: an MF-GLaM (blue curves) and a GLaM trained solely on HF data (red curves), the latter referred to as the “HF-only” GLaM. Results are shown for increasing HF training set sizes, ranging from $N_{\text{H}}=100$ up to $N_{\text{H}}=400$ data points, while the LF training set size remains fixed at $N_{\text{L}}=1,\!000$ data points. We notice that, for all three input points and across all HF training set sizes, the MF-GLaM consistently captures the conditional HF distribution more accurately than the HF-only GLaM. Even for $\boldsymbol{x}_{2}$, where the LF and HF response supports are disjoint, a quite uncommon scenario in practice, the MF‑GLaM reproduces the HF conditional PDF more faithfully compared to relying solely on HF data. We notice that in this four-dimensional application, $N_{\text{H}}=100$ HF training data points are insufficient to achieve an acceptable approximation of the response PDF; however, the MF approach still yields improved predictions compared to the HF-only model. Once the HF training set size increases to $N_{\text{H}}=200$, the MF predictions become very close to the true HF PDF, and with $N_{\text{H}}=400$, the MF approximation is nearly exact. In contrast, the HF-only GLaM exhibits greater variability and struggles to reproduce the PDF shape in certain cases (particularly for inputs $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$).
@@
These observations are further supported by the global approximation performance comparison shown in Fig.˜2, and its numerical summary, provided in Table˜3. The box plot presents the performance comparison in terms of normalized Wasserstein distance $\epsilon_{\text{W}}$ of the MF-GLaM, the HF-only GLaM, and the LF-only GLaM, the latter serving as a baseline. Each box group corresponds to a different HF training set size, ranging from $N_{\text{H}}=100$ to $N_{\text{H}}=800$ points, with $N_{\text{r}}=25$ repetitions per scenario. Throughout these tests, we drew once $N_{\text{r}}$ LF training datasets of size $N_{\text{L}}=1,\!000$ each, which were reused across all $N_{\text{H}}$ values.
We observe that the MF-GLaM consistently outperforms the GLaM trained solely on the HF data and shows faster convergence to the true HF model, indicated by the rapid decrease in Wasserstein distance. Especially for $N_{\text{H}}=200$, the error is reduced by approximately one order of magnitude. The only exception occurs at $N_{\text{H}}=100$, where the median performance of the MF and HF-only emulators appears comparable. This can be attributed to the imposed limitation of the correction term $\lambda_{1}^{\delta,\text{PC}}$ to a polynomial degree of 1 at this training size, introduced to avoid overfitting. However, as shown in Table˜2, the actual discrepancy between the HF and LF expansions of $\lambda_{1}$ involves higher-degree terms. Nonetheless, the MF-GLaM demonstrates reduced variability across repetitions, as indicated by the significantly smaller interquartile range of $\epsilon_{\text{W}}$ in Table˜3, which suggests more robust predictive performance.
Figure 2: Synthetic GLaMs example. Comparison of the normalized Wasserstein distance $\epsilon_{\text{W}}$ for LF-only, HF-only, and MF GLaMs across increasing HF experimental design sizes $N_{\text{H}}$. The box plots correspond to $N_{\text{r}}=25$ repetitions of the procedure, and $\epsilon_{\text{W}}$ is computed using 1,0001,\!000 test points.
Table 3: Synthetic GLaMs example. Median (interquartile range) of the normalized Wasserstein distance $\epsilon_{\text{W}}$ $\times 10^{-2}$, computed over $N_{\text{r}}=25$ repetitions, comparing HF‑only GLaM and MF‑GLaM models for four HF training set sizes. Baseline $\epsilon_{\text{W}}$, for the LF‑only GLaM: $5.8\,(0.1)\times 10^{-2}$.
@@
Finally, we also compare how well the emulators reproduce the mean and variance of the HF stochastic model. Since these two statistical quantities are deterministic, we use a normalized mean‐square error (NMSE) measure to assess their prediction accuracy. Specifically, for each point $\boldsymbol{x}^{(i)}$ in a test set of size $N_{\mathrm{t}}=10^{4}$, the NMSE is given by:
@@
and $\eta_{\text{H}}^{(i)}$ the true value of the mean or variance at $\boldsymbol{x}^{(i)}$ for the HF simulator, given by Eq.˜5 and Eq.˜6, respectively, since in this example our true HF simulator is a GLaM. Also, $\eta_{\text{GLaM}}^{(i)}$ is the corresponding mean or variance value predicted by the GLaM surrogate model.
Fig.˜4(a) shows the NMSE for the mean (LABEL:fig:06Synth_RMSE_mean) and variance (LABEL:fig:06Synth_RMSE_var) estimates of the stochastic emulators whose predictive distributions were evaluated in Fig.˜2. The corresponding box plots align with those for the Wasserstein distance and further confirm that the MF-GLaM provides more accurate and robust predictions compared to its single-fidelity counterparts, also for the statistics of the model response.
(a) Synthetic GLaMs example. Comparison of the normalized mean-squared error $\epsilon_{\text{v}}$ on the prediction of (a) the mean and (b) the variance, for LF-only, HF-only, and MF GLaMs across increasing HF experimental design sizes $N_{\text{H}}$. The box plots correspond to $N_{\text{r}}=25$ repetitions of the procedure, and $\epsilon_{\text{v}}$ is computed using $N_{\text{t}}=10^{4}$ test points.
@@ 4.2 Stochastic borehole function
The input vector $\boldsymbol{x}$ contains eight random variables, whose distributions are shown in Table˜4.
@@
$R_{w}$ [m] $\mathcal{N}\,([0.1,\;0.016])$ radius of borehole
$H_{u}$ [m] $\mathcal{U}\,([990,\;1110])$ potentiometric head of upper aquifer
$K_{w}$ [m/yr] $\mathcal{U}\,([9855,\;12045])$ hydraulic conductivity of borehole
$R$ [m] $\mathcal{LN}\,([7.71,\;1.0056])$ radius of influence
$T_{u}$ $\text{m}^{2}$/yr] $\mathcal{U}\,([63070,\;115600])$ transmissivity of upper aquifer
$T_{l}$ $\text{m}^{2}$/yr] $\mathcal{U}\,([63.1,\;116])$ transmissivity of lower aquifer
$H_{l}$ [m] $\mathcal{U}\,([700,\;820])$ potentiometric head of lower aquifer
$L$ [m] $\mathcal{U}\,([1120,\;1680])$ length of borehole
@@
Moreover, Lüthen et al. (2023) modified the borehole function in Eq.˜38 into a stochastic simulator by treating the five input parameters, $\boldsymbol{\Omega}_{\text{H}}=(R,T_{u},T_{l},H_{l},L)$, as latent variables. We adopt this variation for our application as the HF stochastic simulator, which results in the following three-dimensional HF model: $f_{s,\text{H}}(r_{w},h_{u},k_{w})=f(r_{w},h_{u},k_{w};\omega_{\text{H}})$.
Additionally, we obtain the LF stochastic simulator from the deterministic LF model in Eq.˜39 by considering six of its input parameters to be latent: $\boldsymbol{\Omega}_{\text{L}}=(K_{w},R,T_{u},T_{l},H_{l},L)$. This results in the following two-dimensional LF model: $f_{s,\text{L}}(r_{w},h_{u})=f_{\text{L}}(r_{w},h_{u};\omega_{\text{L}})$.
This example investigates the applicability of the multifidelity GLaM methodology in cases where the LF explicit input variables are a subset of the HF ones. In this example, the two-dimensional LF input is a subset of the three-dimensional HF input. As a result, whereas the two common input variables can be modeled using all the available HF and LF data in the multifidelity GLaM through both the LF and the discrepancy part in Eq.˜19, the additional HF-only input variable and its interactions have to be captured exclusively through the discrepancy component of the MF-GLaM, relying solely on the HF data.
Fig.˜4 compares the PDF approximation capabilities of an MF-GLaM and an HF-only GLaM for three random points. In the same plots, we see the true HF PDF (black dashed curve) as well as the true LF (gray dotted curve) PDF. We notice that the HF-LF relationship can vary across the input space. For example, at $\boldsymbol{x}_{3}$, the LF PDF closely matches the HF reference, whereas at $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$ the LF distribution is more noticeably shifted or differs in shape. The MF-GLaM (blue curve) successfully corrects for these discrepancies, consistently aligning with the true HF response, even when $N_{\text{H}}=100$. In contrast, the HF-only GLaM (red curve) tends to misplace the peak of the PDF or produce much longer tails, especially at smaller training sizes.
@@
Fig.˜5 and its numerical summary in Table˜5 compare the global approximation capabilities of the LF‐only, HF‐only, and MF‐GLaMs in terms of the normalized Wasserstein distance $\epsilon_{\text{W}}$. The size of the HF training set ranges from $N_{\text{H}}=100$ up to $N_{\text{H}}=800$, while the LF training set size is fixed at $N_{\text{L}}=1,\!000$ data points. Overall, the MF‐GLaM (blue boxes) consistently outperforms the HF‐only GLaM (red boxes), especially when the HF training set is small. Particularly, with only $N_{\text{H}}=100$ high‐fidelity samples, the MF-GLaM achieves accuracy comparable to that of the HF‐only model trained with $N_{\text{H}}=400$. As $N_{\text{H}}$ increases to 800, the gap between the MF and HF‐only approaches narrows; however, such a large high‐fidelity to low-fidelity data size ratio is generally atypical for multi‐fidelity scenarios.
Figure 5: Stochastic borehole example. Comparison of the normalized Wasserstein distance $\epsilon_{\text{W}}$ for LF-only, HF-only, and MF GLaMs across increasing HF experimental design sizes $N_{\text{H}}$. The box plots correspond to 25 repetitions of the procedure, and $\epsilon_{\text{W}}$ is computed using 1,0001,\!000 test points.
Table 5: Stochastic borehole example. Median (interquartile range) of the normalized Wasserstein distance $\epsilon_{\text{W}}$ $\times 10^{-2}$, computed over $N_{\text{r}}=25$ repetitions, comparing HF‑only GLaM and MF‑GLaM models for four HF training set sizes. Baseline $\epsilon_{\text{W}}$, for the LF‑only GLaM: $2.9\,(0.2)\times 10^{-2}$.